OpenStudy (anonymous):

Write a point-slope equation for the line that has slope 3 and passes through the point (5, 21). Do not use parenthesis on the y side.

OpenStudy (sheraz12345):

USe equation y-y1=m(x-x1) Substitute the values

OpenStudy (anonymous):

y-21=3x-15 y=3x+6